# Generalized Short Path Algorithms: Towards Super-Quadratic Speedup over Markov Chain Search for Combinatorial Optimization

### v1

@@ 3.2 Applications
The fact that our main results in Theorems 3.3 and 3.4 rely on functional inequalities implies many possibilities for interesting algorithmic speedups. Once a Markov Chain with the right properties (log-Sobolev or Poincaré with the proper parameters) is identified, we can derive conditions on cost functions that for which we have super-quadratic speedup over Markov Chains. As a simple toy example, consider an expander graph of size $2^{n}$, where we are given access to a polynomial time oracle that outputs the edges incident on any vertex. Since the graph is an expander, the graph random walk satisfies a Poincaré inequality with constant $\delta$. Consider any assignment of costs to the nodes such that the difference in cost between any two endpoints of an edge is bounded above by a constant. It follows from our results, that the generalized short path framework can find the node with minimum cost with super-quadratically fewer queries than searching with the random walk on the graph. A systematic study of cost functions that yield a speedup for various Markov Chains may lead to some interesting insights. For this paper, however, we focus on identifying connections to problems of general interest.
@@ 3.2.1 Optimization with Fixed Hamming Weight
In Section 5.1 we prove that the Generalized Short Path Framework achieves an overall runtime of $\mathcal{O}^{\ast}\left(\binom{n}{k}^{0.5-c}\right)$ for MaxCut-Hamming on Erdős-Rényi random graphs when $k=\Theta(n)$ (which includes MaxBisection as a special case). We also demonstrate that under the assumption of spectral density, our approach achieves a runtime of $\mathcal{O}^{\ast}\left(\binom{n}{k}^{0.5-c(n)}\right)$ with $c(n)=\Theta\left(\log(n/k)^{-1}\right)$, and thus the super-quadratic advantage over Markov Chain search decays as $n\to\infty$ (similar to the results of  [Has18a]).
@@ 3.2.2 Glauber Dynamics
Glauber Dynamics [Gla63] is a well known sampling algorithm designed to sample from the Gibbs measures corresponding to Hamiltonians such as the Ising or Hardcore models. Since sampling from a Gibbs measure at arbitrarily high inverse temperatures is equivalent to exact optimization, for most hard problems there exists a critical threshold beyond which the Glauber dynamics no longer mixes efficiently. Performing Markov Chain search with Glauber dynamics has two advantages, if the problem is constrained then it provides a natural way to search with a distribution whose support is restricted to feasible solutions only. On the other hand, if the Glauber dynamics mixes for positive inverse temperatures, then low-energy solutions are disproportionately favored compared to the uniform distribution and the result in Markov Chain Search is asymptotically faster than unstructured search (see Lemma 5.20. In each case, we will consider classical algorithms that search using the Glauber dynamics at an inverse temperature slightly below the critical threshold where mixing takes exponential time. We demonstrate in Section 5.2 that for three models of interest, we obtain super-quadratic speedups over polynomially-mixing Glauber dynamics. These models are:
@@
    The Maximum Independent Set problem (or hardcore model) on graphs of constant maximum degree: For this problem the Glauber dynamics is shown to mix only upto critical temperatures that are negative. This means that our starting distribution favors small sets compared to large sets. However, there is the advantage that the Gibbs distribution has support only on independent sets (which are usually much fewer in number than $2^{n}$, which is the total number of subsets). In the case of random regular graphs of sufficiently high degree, we show that this Markov Chain Search algorithm is faster than unstructured search as well as the best known combinatorial algorithms for Maximum Independent Set.
@@
    The Ising Model on random regular graphs of constant maximum degree: The Ising Model is an unconstrained optimization problem and there are $2^{n}$ feasible solutions. In this setting, the Glauber dynamics mixes upto a positive critical inverse-temperature, thus the starting stationary distribution favors low energy solutions and the Markov Chain search is faster than unstructured search.
@@
    The Sherrington Kirkpatrick Model: Like the Ising Model, the Sherringtok Kirkpatrick model is also unconstrained and the Glauber dynamics mixes upto a positive inverse temperature. In this case, however, the exponent of our advantage over quadratic speedup follows with $n$. Specifically, we show a quantum runtime of $\mathcal{O}^{\ast}\left((\pi(E^{\ast})^{0.5-c(n)}\right)$ where $c(n)=\Theta(1/\log(n))$.
3.2.3 Super-Quadratic Speedup Over any Polynomial Time Gibbs Sampler
Our result regarding the Maximum Independent Set (or, hardcore model) on graphs of bounded degree allows us to go a step further and argue that it is very likely that the generalized short path can achieve a super-Grover speedup over all polynomial time Gibbs samplers (whether or not they are based on Gibbs sampling). The key observation is that for the hardcore model there is a critical fugacity $\lambda_{c}(d)=\exp(-\beta_{c}(d))=\frac{(d-1)^{d-1}}{(d-2)^{d-2}}$ such that for graphs of maximum degree $d$, Glauber dynamics mixes in time $\mathcal{O}(n\log(n))$ for any $\lambda<\lambda_{c}(d)$. However, it has been shown that computing the partition function of the hardcore model is NP-hard for any fugacity $\lambda>\lambda_{c}(d)$ [Sly10, SS14]. Due to the well known reduction between sampling and counting, it must therefore be NP-hard to sample from the corresponding Gibbs distribution. We therefore have an almost complete classification of the Gibbs measures corresponding to the hardcore model at different fugacity: it is either NP-hard to sample from the distribution, or the Glauber dynamics mixes in polynomial time. Our results in Section 5.2.1 establish that in the latter setting, the “short jump” of Algorithm 2 takes only polynomial time and produces a state $|\psi_{b}\rangle$ such that it is NP-hard to sample from any Gibbs distribution $\pi_{\beta}$ obeying $\pi_{\beta}(z^{\ast})=\Omega\left(|\langle\psi_{b}|z^{\ast}\rangle|^{2}\right)$.
Consequently, there exists a quantitative separation between the ground states of the short path Hamiltonian and Gibbs distributions. It is notable that the Markov Chain methods are obstructed at the critical fugacity due to the development of long range correlations. Such long range correlations in the ground state are also likely to lead to a vanishing gap for the short path Hamiltonian. This indicates that the ground state of the short path algorithm increases the overlap with the ground state, without creating long range correlations. Further understanding of the qualitative differences between the ground state and classical measures such as the Gibbs distribution, may shed light on the mechanisms of the short path algorithm.
We note that we resort to computational assumptions above only to show that the ground state $|\psi_{b}\rangle$ has higher overlap with the optimal solution than an efficiently sampleable Gibbs distribution. It follows unconditionally from our analysis that the ground state itself does not encode a Gibbs distribution beyond the critical mixing threshold (or indeed any distribution using which we can sample from such a Gibbs distribution by rejection sampling). In particular, Lemma 4.20 shows that for any ground state prepared by a short jump that is covered by our analysis, the trace distance from the starting state, and hence the total variation distance from the starting distribution is exponentially small.
@@ 4 Generalized Short Path Framework
This section provides a simplified and generalized analysis of the short path algorithm presented in [DPCB23]. It also highlights limitations of the current method of analysis, and describes a general recipe for the determining a speedup over Grover.
@@ 4.1 Summary of main results
In what follows, let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible Markov chain over a finite set $\mathcal{X}$. Here $\pi$ denotes a stationary distribution that is uniform over $\mathcal{X}$ and $P$. It is assumed that the spectral gap of $P$ is lower-bounded by $\delta$. Using the discriminant matrix of $P$, we can define a more general short-path Hamiltonian $H_{b}$ that allows one to work with mixing operators other than $-\frac{X}{n}$.
Definition 4.1 (Short-path Hamiltonian $H_{b}$).
Consider a reversible Markov chain $\mathcal{M}=(\mathcal{X},P,\pi)$. Let $H:\mathcal{X}\mapsto\mathbb{R}$ be a cost Hamiltonian. The short-path Hamiltonian $H_{b}$ is given by
@@
More generally $g_{\eta}:[-1,\infty)\mapsto[-1,0]$ can be a non-decreasing, concave function that is differentiable at every point where it is non-zero. However, the specific choice we make is sufficient for our purposes. We will also sometimes refer use the notation $G_{\eta}:=g_{\eta}\left(\frac{H}{|E^{\ast}|}\right)$.
One major component of the analysis of Algorithm 2 is determining an upper bound on $b$ for which the spectral gap of the short-path Hamiltonian $H_{b}$ is still large, i.e., $\Omega\left(\frac{1}{\operatorname{poly}(n)}\right)$. This upper bound serves as a proxy for how large the “short jump” is. A second major component is determining the increased overlap with the optimal solution provided by the short-jump. To do so, we rely on the definition of $\Delta_{P}$ stability, which we restate below.
Definition 4.2 ($\Delta_{P}$ stability).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain. We say that the cost Hamiltonian $H$ is $\Delta_{P}(\eta)$ stable under $\mathcal{M}$ if
[…]
where $h_{\eta}(x):=g_{\eta}(\frac{x}{\lvert E^{*}\rvert})$.
@@
If the short jump can be accomplished efficiently, then $\Delta_{P}$ stability captures whether the short path approach provides a super-Grover runtime. This condition also has an intuitive interpretation. If we consider the optimization landscape defined by $\mathcal{M}$, $H$ and a well (controlled by $\eta$) around the global minimum with energy $E^{*}$, then we do not want the energy to increase too much when moving within and around the well. Specifically, for a super-Grover runtime it should hold that $\Theta\left(\frac{\lvert E^{*}\rvert}{\ln(1/\pi(E^{*}))}\right)$, where $\pi$ is the stationary distribution of $\mathcal{M}$. It is worth remarking that if $\eta=0$, we recover the quantum unstructured search algorithm.
@@
It turns out that any upper bound on $\Delta_{P}(\eta)$ suffices when bounding the runtime. For example, it is a simple consequence of Jensen’s inequality that we can take $\Delta_{P}(\eta)$ to be $\sqrt{\lVert\psi\rVert_{P}}$ with $\psi=H$. A key technical contribution of this work is to reduce the conditions for determining whether a super-Grover runtime is possible to determining the log-Sobolev constant $\omega$ and the $P$-pseudo Lipschitz norm $\lVert H\rVert_{P}$ (or $\Delta_{P}$) for cost Hamiltonian $H$.
We summarize our main result below:
Theorem 4.3.
@@
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible Markov chain, and let $H:\mathcal{X}\mapsto\mathbb{R}$ be a diagonal, $\Delta_{P}$-stable Hamiltonian with ground state energy $E^{\ast}$, $P$-pseudo Lipschitz norm $\lVert H\rVert_{P}$. In addition, suppose $\mathcal{M}$ has a log-Sobolev constant $\omega$. If $b$ satisfies
$b<b^{\ast}:=\frac{2}{3}\gamma\omega\ln\left(\frac{1}{\pi(E^{\ast})}\right),$
where
@@
$\gamma=\frac{\omega((1-\eta)E^{\ast}-\mathbb{E}_{\pi}[H])^{2}}{\lVert H\rVert_{P}\ln(1/\pi(E^{\ast}))},$
@@
then there exists a short-path algorithm that determines the ground state of $H$ over $\mathcal{X}$ with running time
@@
The proof is evident after combining the statements of Theorems 4.28, 4.15, Lemma 4.25, and Corollary 4.9
@@
Theorem 4.4.
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible Markov chain, and let $H:\mathcal{X}\mapsto\mathbb{R}$ be a diagonal, $\Delta_{P}$-stable Hamiltonian with ground state energy $E^{\ast}$, $P$-pseudo Lipschitz norm $\lVert H\rVert_{P}$. In addition, suppose $\mathcal{M}$ has a Poincaré constant $\delta$. If $b$ satisfies
$b<b^{\ast}:=\delta\frac{4\sqrt{6}-1}{10},$
then there exists a short-path algorithm that determines the ground state of $H$ over $\mathcal{X}$ with running time
@@
The proof is evident after combining the statements of Theorems 4.28, 4.17, Lemma 4.25, and Corollary 4.10. ∎
In general, the log-Sobolev constant $\omega$ can be significantly smaller than the spectral gap of the chain $\delta$, however, we argue that this is not the case when Theorem 4.3 provides a super-Grover runtime. Specifically, Theorem 4.3 requires that $b^{*}$ is a constant, and by extension, implies we need $\omega^{-1}=\Theta(\ln(1/\pi(E^{*})))$. For example, for a very hard problem, where Markov Chain search finds an optimal assignment with exponentially-small probability, i.e., $\Theta(\ln(1/\pi(E^{*})))=\Theta(n)$, the condition on $b^{*}$ will imply that $\omega$ will be large, i.e., $\Omega\left(\frac{1}{\operatorname{poly}(n)}\right)$. Thus, in cases where it provides a super-Grover runtime, Theorem 4.3 asserts that we do not get a slower runtime by using $\omega$ instead of $\delta$.
We have the following evident corollary of the above results.
Corollary 4.5.
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible Markov chain, and let $H:\mathcal{X}\mapsto\mathbb{R}$ be a diagonal Hamiltonian with ground state energy $E^{\ast}$. Suppose the pair $(\mathcal{M},H)$ result in $b^{\ast}$ and $\gamma$ that are independent of $n$. If
$\frac{\lvert E^{\ast}\rvert}{\Delta_{P}}=\Theta\left(\ln\left(1/\pi(E^{\ast})\right)\right),$
then under Assumption 1 there exists a short-path algorithm that finds the minimizer of $H$ over $\mathcal{X}$ with super-Grover running time. The advantage over Grover is:
$\frac{\eta(1-\eta)\lvert E^{\ast}\rvert b^{\ast}}{\Delta_{P}}.$

### v2

@@ 3.2 Applications
The fact that our main results in Theorems 3.3 and 3.4 rely on functional inequalities implies many possibilities for interesting algorithmic speedups. Once a Markov chain with the right properties (log-Sobolev or Poincaré with the proper parameters) is identified, we can derive conditions on cost functions for which we have super-quadratic speedup over Markov chains. As a simple toy example, consider an expander graph of size $2^{n}$, where we are given access to a polynomial time oracle that outputs the edges incident on any vertex. Since the graph is an expander, the graph random walk satisfies a Poincaré inequality with constant $\delta$. Consider any assignment of costs to the nodes such that the difference in cost between any two endpoints of an edge is bounded above by a constant $\Delta$. It follows from our results (Theorem 3.4), that the generalized short path framework can find the node with minimum cost with super-quadratically fewer queries than searching with the random walk on the graph. A systematic study of cost functions that yield a speedup for various Markov chains may lead to some interesting insights. For this paper, however, we focus on identifying connections to problems of general interest.
Theorem 3.5 presents a summary of the quantum runtimes obtained for all applications considered in this paper.
Theorem 3.5 (Applications of Generalized Short Path Framework).
The following optimization problems exhibit a quantum runtime of $\mathcal{O}^{\ast}\left(T(n)^{0.5-c(n)}\right)$, where $T(n)$ is the runtime of a classical algorithm based on Markov chain search, and $c(n)>0$, specified in Tables 1 and 2, quantifies the degree of improvement over quadratic speedup:
  1. 1.
    Finding a generalization of maximum bisection (called MaxCut Hamming) over Erdős-Rényi random graphs with constant average degree, where the smaller partition is constrained to have $k$ nodes (where $k\leq n/2$). In both cases, the Markov chain is the transposition walk.
  2. 2.
    Finding a maximum independent set on graphs with constant maximum degree, where the Markov chain is Glauber dynamics,
  3. 3.
    Optimizing the antiferromagnetic Ising model on graphs of constant maximum degree, where the Markov chain is Glauber dynamics
  4. 4.
    Optimizing the Sherrington-Kirkpatrick model where the Markov chain is the Glauber dynamics
Problem Graph Model Advantage $(c(n))$ Theorem
Maximum Bisection $\mathcal{G}_{\textup{ER}}(n,\frac{d}{n}),d=\mathcal{O}_{n}(1)$ $\Theta(1)$ Theorem 5.6
MaxCut Hamming $\mathcal{G}_{\textup{ER}}(n,\frac{d}{n}),d=\mathcal{O}_{n}(1)$ $\Theta(1/\log(n/k))$ Theorem 5.6
Table 1: Degree of quantum improvement for problems using MCS with transposition walk.
Model Graph Model Inverse Temperature Advantage $(c(n))$ Theorem
Hardcore Model $d=\mathcal{O}_{n}(1)$ $\beta\geq-(d-1)\ln(d-1)+d\ln(d-2)$ $\Theta(1)$ Theorem 5.17
$d=\mathcal{O}_{n}(1)$ $\beta=-(d-1)\ln(d-1)+d\ln(d-2)$ $\Omega\left(\frac{1}{n^{1+4e(1+\frac{1}{d-2})}}\right)$ Theorem 5.19
$\mathcal{G}_{\textup{RG}}(n,d),d=\mathcal{O}_{n}(1)$ $\beta>\ln\left({2\sqrt{d-1}-1}\right)$ $\Omega(1/n)$ Theorem 5.21
Anti- ferromagnetic Ising Model $d=\mathcal{O}_{n}(1)$ $(d-1)\tanh(\beta)<1$ $\Theta(1)$ Theorem 5.26
$d=\mathcal{O}_{n}(1)$ $(d-1)\tanh(\beta)\leq 1+o_{n}(1)$ $\Omega(1/n^{1+1/d})$ Theorem 5.28
$\mathcal{G}_{\textup{RG}}(n,d),d=\mathcal{O}_{n}(1)$ $(d-1)\tanh(\beta)\leq\frac{d-1}{8\sqrt{d-1}-1}$ $\Theta(1)$ Theorem 5.31
Sherrington-Kirkpatrick Model $\beta<\frac{1}{4}$ $\Theta(1/\log(n))$ Theorem 5.35
Table 2: Degree of quantum improvement for problems using MCS with Glauber Dynamics. Note in all cases $d\geq 3$. $\beta$ corresponds to inverse temperature for the Gibbs sample $\propto\exp(-\beta H)$.
Quantum Improvements that Lead to Speedups over Classical State-of-the-art
We would like to highlight two particular cases where we obtain quantum algorithms that are faster than existing provable, classical state-of-the-art generic algorithms. The first case is for the antiferromagnetic Ising model over random regular graphs. Here in Table 2 we obtain $c(n)=\Theta(1)$, leading to an asymptotic super-quadratic speedup over Markoc chain search. The underlying Markov chain is the state-of-the-art classical approach of [CCC+25]. Secondly, while for the case of MIS over random-regular graphs the improvement over quadratic speedup falls with $n$, the underlying classical MCS algorithm appears to be faster than other classical generic approaches when the degree is sufficiently large.
In the following subsections, we expand more on the details of the applications that we consider and how the ability to speedup MCS enables faster quantum algorithms. In Section 3.2.3, we also discuss about potential evidence that some of the speedups obtained cannot be what we call “Groverized.” We say that a speedup with (generalized) short-path can be Groverized if the edge-over-quadratic can be replicated with a classical subroutine combined with quantum minimum finding.
@@ 3.2.1 Optimization with Fixed Hamming Weight
The runtimes we obtain are highlighted in Table 1. In Section 5.1 we prove that the Generalized Short Path Framework achieves an overall runtime of $\mathcal{O}^{\ast}\left(\binom{n}{k}^{0.5-c}\right)$ for MaxCut-Hamming on Erdős-Rényi random graphs when $k=\Theta(n)$ (which includes MaxBisection as a special case). We also demonstrate that under the assumption of spectral density, our approach achieves a runtime of $\mathcal{O}^{\ast}\left(\binom{n}{k}^{0.5-c(n)}\right)$ with $c(n)=\Theta\left([\log(n/k)]^{-1}\right)$, when $k=o(n)$, and thus the super-quadratic advantage over Markov chain search decays as $n\to\infty$ (similar to the results of  [Has18a]).
@@ 3.2.2 Glauber Dynamics
Glauber dynamics [Gla63] is a well known sampling algorithm designed to sample from the Gibbs measures corresponding to Hamiltonians such as the Ising or Hardcore models. Since sampling from a Gibbs measure at arbitrarily high inverse temperatures solves the exact optimization, for most hard problems there exists a critical threshold beyond which the Glauber dynamics no longer mixes efficiently, which is discussed in Section 3.2.3. Performing Markov chain search with Glauber dynamics has two advantages, if the problem is constrained then it provides a natural way to search with a distribution whose support is restricted to feasible solutions only. On the other hand, if the Glauber dynamics mixes for positive inverse temperatures, then low-energy solutions are more favored compared to the uniform distribution and the result in Markov chain search is asymptotically faster than unstructured search. In each case, we will consider classical search algorithms that use the Glauber dynamics at for inverse-temperatures in the ranges specified in Table 2. We demonstrate in Section 5.2 that for three models of interest, we obtain super-quadratic, but potentially asymptotically falling, speedups over polynomially-mixing Glauber dynamics. These models are:
@@
    The Hardcore model on graphs of constant maximum degree: For this problem the Glauber dynamics is shown to mix only up to critical temperatures that are negative. This means that our starting distribution favors small sets compared to large sets. However, there is the advantage that the Gibbs distribution has support only on independent sets (which are usually much fewer in number than $2^{n}$, which is the total number of subsets). In the case of random regular graphs of sufficiently high degree, we show that this Markov chain search algorithm is faster than unstructured search (Proposition 5.22) as well as the best known combinatorial algorithms for finding a maximum independent set.
@@
    The Antiferromagnetic Ising model on graphs of constant maximum degree: The Antiferromagnetic Ising model is an unconstrained optimization problem, encompassing problems like MaxCut, and there are $2^{n}$ feasible solutions. In this setting, the Glauber dynamics mixes up to a positive critical inverse-temperature, thus the starting stationary distribution favors low energy solutions and the Markov chain search is faster than unstructured search (Proposition 5.32).
@@
    The Sherrington-Kirkpatrick model: Like the Ising model, the Sherrington-Kirkpatrick model is also unconstrained. The Glauber dynamics has been shown to have fast mixing in total variation distance up to inverse-temperature $\beta<\frac{1}{4}$ [CE22]. We show that the quantum short-path algorithm applies to Glauber dynamics for SK within this $\beta$ range. Additionally MCS for this $\beta$ range outperforms bruteforce search (Proposition 5.36). However, mixing in Wasserstein-2 is possible all the way up to a known computational threshold of $\beta<1$ [AMS24, Cel22].
3.2.3 Can the Generalized Short-Path Algorithm be “Groverized”?
Our techniques also allow us to shed more light on a fundamental question about the viability of true super-quadratic speedups with the short path framework. The algorithms in this paper as well as the earlier frameworks, rely on preparing a quantum state whose overlap with the global minimizers is larger than that of some easily prepared starting state, referred to as the short-jump, and then applying quantum minimum finding to that state. It is apparent that if there existed a classical algorithm to sample in polynomial time from a distribution with overlap that matches that of this intermediate state, then there is a classical algorithm that finds the global minimum only quadratically slower than the short path algorithm. The advantage over search is then essentially “Groverized” and there is no hope for true super-quadratic speedups. It is therefore important to understand to what degree classical sampling techniques can approach the overlap of the intermediate state, as discussed in [DPCB23]. Most natural sampling algorithms are based on the analysis of Markov chains and so our framework provides a useful tool to probe this question.
The quantum speedups over MCS with Gibbs samplers, e.g., Glauber dynamics, are potentially interesting candidates for providing evidence against Groverization. This is because there are known computational phase transitions in mixing time. Particularly there are two known thresholds for two different graph models. For constant-degree graphs, this is known as the tree-uniqueness threshold [Sly10, GSV16] and for sparse, random-regular graphs (and the SK model) this known as the onset of disorder chaos [Cha09, AMS24, HPP24]. The first threshold holds unless $\mathsf{NP}=\mathsf{RP},$ and the latter holds against any “stable” Gibbs sampling algorithm [AMS24]. If the quantum algorithm can produce an overlap with the optimal solution that is larger than the Gibbs distribution at either of these thresholds, then this is evidence against dequantization. This is because, in most cases, Gibbs sampling is the only “natural” classical thing to do.
For the case of bounded-degree graphs, we find that the improvement over quadratic, i.e. benefit from the intermediate state, falls with $n$ at the uniqueness theshold. However, this may only be a limitation of current proof techniques. Our numerical results show that the conditions for asymptotic super-quadratic speedup with the generalized short path framework tend to be overly pessimistic.
For the random $d$-regular case, it appears unclear if Glauber dynamics for the antiferromagnetic Ising model can mix all-the-way to the point of disorder chaos. However, the current state-of-the-art inverse temperature $\beta_{T}$ [CCC+25] is close, i.e. both $\beta_{T}$ and the point of disorder chaos $\beta_{\textup{DC}}$ scale inversely with $\sqrt{d}$. Additionally, we prove that the generalized short-path algorithm attains an asymptotic super-quadratic speedup at $\beta_{T}$.
Given the above two points, there is still hope that such evidence against Groverization can be shown.
Remark 3.6.
An earlier version of the paper only included quantum speedups over Glaubers dynamics for inverse temperatures that are a constant fraction away from the tree-uniqueness threshold. Recent results in the classical literature have shown that Glauber dynamics can mix up to the uniqueness threshold and beyond the threshold when the ensemble is random-regular graphs. As highlighted above, this has led to some new quantum runtimes along with some reassessments of the original claims against Groverization. Specifically, Glauber dynamics has been shown to mix for larger inverse-temperatures than previously thought.
@@ 4 Generalized Short Path Framework
This section provides a simplified and generalized analysis of the short path algorithm presented in [DPCB23]. It also highlights limitations of the current method of analysis, and describes a general recipe for the determining a super-quadratic speedup.
@@ 4.1 Summary of main results
In what follows, let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible, aperiodic [^5] Note that any Markov chain can be made aperiodic by making it lazy: $P\mapsto\frac{I+P}{2},D(P)\mapsto\frac{D(P)+I}{2}$. We will generally assume that this transformation has been implicitly done if the chain is not aperiodic, as the assumption only acts to simplify some results and is not critical. Markov chain over a finite set $\mathcal{X}$. Here $\pi$ denotes the stationary distribution of $P$. It is assumed that the spectral gap of $P$ is lower-bounded by $\delta$. Using the discriminant matrix of $P$, we can define a more general short path Hamiltonian $H_{b}$ that allows one to work with mixing operators other than $\frac{X}{n}$.
Definition 4.1 (Generalized Short Path Hamiltonian $H_{b}$).
Consider a reversible Markov chain $\mathcal{M}=(\mathcal{X},P,\pi)$. Let $H:\mathcal{X}\rightarrow\mathbb{R}$ be a cost Hamiltonian with $E^{\ast}<0$. The short path Hamiltonian $H_{b}$ is given by
@@
More generally $g_{\eta}:[-1,\infty)\rightarrow[-1,0]$ can be a non-decreasing, concave function that is differentiable at every point where it is non-zero. However, the specific choice we make is sufficient for our purposes. We will also sometimes use the notation $G_{\eta}:=g_{\eta}\left(\frac{H}{|E^{\ast}|}\right)$.
@@
One major component of the analysis of Algorithm 2 is determining an upper bound on $b$ for which the spectral gap of the short path Hamiltonian $H_{b}$ is still large, i.e., $\Omega\left(\frac{1}{\operatorname{poly}(n)}\right)$. This upper bound serves as a proxy for how large the “short jump” is. A second major component is determining the increased overlap with the optimal solution provided by the short jump. To do so, we rely on the definition of $\Delta_{P}$ stability (Definition 3.1). If the short jump can be accomplished efficiently for some constant $b$, then $\Delta_{P}$ stability captures whether the short path approach provides a super-Grover runtime. This condition also has an intuitive interpretation. If we consider the optimization landscape defined by $\mathcal{M}$, $H$ and a well (controlled by $\eta$) around the global minimum with energy $E^{*}$, then we do not want the energy to increase too much when moving within and around the well. Specifically, for a super-Grover runtime it should hold that $\Delta_{P}(\eta)=\Theta\left(\frac{\lvert E^{*}\rvert}{\ln(1/\pi(E^{*}))}\right)$ for some $\eta$, where $\pi$ is the stationary distribution of $\mathcal{M}$. It is worth remarking that if $\eta=0$, we recover the quantum unstructured search algorithm. When the explicit value of $\eta$ such that Definition 3.1 holds is not important, we will simply denote $\Delta_{P}(\eta)$ as $\Delta_{P}$.
@@
It turns out that any upper bound on $\Delta_{P}$ suffices when bounding the runtime. For example, it is a simple consequence of Jensen’s inequality that we can take $\Delta_{P}$ to be $\sqrt{\lVert\psi\rVert_{P}}$ with $\psi=H$. A key technical contribution of this work is to reduce the conditions for determining whether a super-Grover runtime is possible to determining the log-Sobolev constant $\omega$, or Poincaré constant $\delta$, and the $P$-pseudo Lipschitz norm $\lVert H\rVert_{P}$ for cost Hamiltonian $H$.
We summarize our main results below, which are the formal versions of Theorems 3.3 and 3.4:
Theorem 4.2 (Short-Path Runtime Under log-Sobolev).
@@
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible, aperiodic Markov chain, and let $H:\mathcal{X}\rightarrow\mathbb{R}$ be a diagonal, $\Delta_{P}$-stable Hamiltonian (Definition 3.1) with ground state energy $E^{\ast}<0$, $P$-pseudo Lipschitz norm $\lVert H\rVert_{P}$. In addition, suppose $\mathcal{M}$ has a log-Sobolev constant lower bounded by $\omega$. Then $\gamma$-spectral density (Definition 3.2) is satisfied for
Also if $b$ satisfies
$b<b^{\ast}:=\frac{2}{3}\gamma\omega\ln\left(\frac{1}{\pi(E^{\ast})}\right)$
and
@@
$\frac{1}{3\gamma\ln(1/\pi(E^{\ast}))\sqrt{\mathbb{P}_{\pi}(E\leq(1-\eta)E^{*})}}-\max\left(\frac{4\ln(\lvert\mathcal{X}\rvert)}{\omega},\frac{3\lvert E^{\ast}\rvert^{2}(1-\eta)^{2}}{\Delta_{P}^{2}}\right)\geq 2,$ (11)
@@
then there exists a short path algorithm that determines the ground state of $H$ over $\mathcal{X}$ with running time
@@
The proof is evident after combining the statements of Corollary 4.22, Theorem 4.14, Lemma 4.18, Lemma 4.10, and Corollary 4.8
@@
Theorem 4.3 (Short-Path Runtime Under Poincaré).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible, aperiodic Markov chain, and let $H:\mathcal{X}\rightarrow\mathbb{R}$ be a diagonal, $\Delta_{P}$-stable Hamiltonian (Definition 3.1) with ground state energy $E^{\ast}<0$, $P$-pseudo Lipschitz norm $\lVert H\rVert_{P}$. In addition, suppose $\mathcal{M}$ has a Poincaré constant $\delta$. Then $\gamma$-spectral density (Definition 3.2) is satisfied for
$\gamma=\frac{\sqrt{\delta}((1-\eta)E^{\ast}-\mathbb{E}_{\pi}[H])}{\sqrt{\lVert H\rVert_{P}}\ln(1/\pi(E^{\ast}))},$
and suppose that $\frac{\delta}{2}>\pi(E^{\ast})^{\gamma}$. Also if $b$ satisfies
$b<b^{\ast}:=\frac{\delta}{4}$
and
$\frac{1}{2\sqrt{\mathbb{P}_{\pi}(E\leq(1-\eta)E^{*})}}-\max\left(\frac{4\ln(\lvert\mathcal{X}\rvert)}{\delta},\frac{3\lvert E^{\ast}\rvert^{2}(1-\eta)^{2}}{\Delta_{P}^{2}}\right)\geq 2,$ (12)
then there exists a short path algorithm that determines the ground state of $H$ over $\mathcal{X}$ with running time
@@
The proof is evident after combining the statements of Corollary 4.22, Theorem 4.16, Lemma 4.10, Lemma 4.18, and Corollary 4.9. ∎
Remark 4.4.
The conditions in (11) and (12) are more of a technical nature and actually very weak. Specifically, the two quantities that we are subtracting will likely be separated by exponential factors in $n$, e.g. $\mathbb{P}_{\pi}(E\leq(1-\eta)E^{*})$ will be exponentially small for hard problems. This is discussed in Section 4.2.2. For similar reasons the additional condition on $\delta,\pi(E^{\ast}),\gamma$ in Theorem 4.3 is also technical and very weak. This is discussed in Section 4.2.1.
In general, the log-Sobolev constant $\omega$ can be significantly smaller than the spectral gap of the chain $\delta$, however, we argue that this is not the case when Theorem 4.2 provides a super-Grover runtime. Specifically, Theorem 4.2 requires that $b$ is a constant, and by extension, implies we need $\omega^{-1}=\Theta(\ln(1/\pi(E^{*})))$. For example, for a very hard problem, where Markov chain search finds an optimal assignment with exponentially-small probability, i.e., $\Theta(\ln(1/\pi(E^{*})))=\Theta(n)$, the condition on $b$ will imply that $\omega$ will be large, i.e., $\Omega\left(\frac{1}{\operatorname{poly}(n)}\right)$. Thus, in cases where it provides a super-Grover runtime, Theorem 4.2 asserts that we do not get a slower runtime by using $\omega$ instead of $\delta$.
We have the following evident corollary of the above results. This provides conditions under which the advantage over Grover with short-path does not asymptotically fall with the problem size.
Corollary 4.5 (Conditions for Asymptotic Super-Grover Runtime).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible, aperiodic Markov chain, and let $H:\mathcal{X}\rightarrow\mathbb{R}$ be a diagonal Hamiltonian with ground state energy $E^{\ast}<0$. Suppose that the chain $\mathcal{M}$ and cost-function $H$ are such that either Theorem 4.2 or 4.3 is satisfied for some $b=\Theta(1)$ and
$\frac{\lvert E^{\ast}\rvert}{\Delta_{P}}=\Theta\left(\ln\left(1/\pi(E^{\ast})\right)\right).$
Then under Assumption 1 the short path algorithm finds the minimizer of $H$ over $\mathcal{X}$ with running-time bounded by
$\mathcal{O}^{\ast}\left([\pi(E^{\ast})^{-1}]^{\frac{1}{2}-c}\right),$
where
$c=\frac{\eta(1-\eta)\lvert E^{\ast}\rvert b}{2\Delta_{P}\ln(1/\pi(E^{\ast}))}=\Theta(1).$